OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 18 ≤10 and x + 18≥ 5

OpenStudy (saifoo.khan):

@TuringTest

OpenStudy (saifoo.khan):

@robtobey

OpenStudy (anonymous):

What is the problem @saifoo.khan can't you help him??

OpenStudy (saifoo.khan):

@waterineyes , im not sure.. i might get this wrong.

OpenStudy (anonymous):

I am also searching for the options.. Where are the following options @googgsh

OpenStudy (anonymous):

Are you having options with you @googgsh

OpenStudy (anonymous):

\[x + 18 \le 10 \implies x \le -8\] \[x + 18 \ge 5 \implies x \ge -13\]

OpenStudy (anonymous):

Can you give the options to me or not @googgsh

OpenStudy (anonymous):

As your wish.. Solve it yourself. And please reply when someone asks you..